# An Agentic AI Workflow to Simplify Parameter Estimation of Complex Differential Equation Systems

### v2

@@ Abstract
Parameter identification for mechanistic Ordinary Differential Equation (ODE) models underpins prediction and control in several applications, yet remains a labor-intensive and brittle process: datasets are noisy and partial, models can be stiff or misspecified, and differentiable implementations demand framework expertise. An agentic AI workflow is presented that converts a lightweight, human-readable specification into a compiled, parallel, and differentiable calibration pipeline. Users supply an XML description of the problem and fill in a Python code skeleton; the agent automatically validates consistency between spec and code, and auto-remediates common pathologies. It transforms Python callables into pure JAX functions for efficient just-in-time compilation and parallelization. The system then orchestrates a two-stage search comprising global exploration of the parameter space followed by gradient-based refinement. The result is an AD-native, reproducible workflow that lowers the barrier to advanced calibration while preserving expert control. An open-source implementation with a documented API and examples is released, enabling rapid movement from problem statement to interpretable ODE models with minimal effort.
@@ 1 Introduction
Ordinary differential equation (ODE) parameter fitting is a foundational task for turning mechanistic models or ideas into predictive tools across the sciences and engineering. In electrochemistry and battery systems, ODEs govern degradation kinetics, thermal behavior, and equivalent-circuit dynamics [1, 2]; accurate parameters are essential for state estimation, control, and safety analysis. In combustion modeling, reactions are often represented using Arrhenius ODE systems [3, 4]. In biology, systems of ODEs describe gene regulation, signaling, and epidemiological dynamics [5]. More broadly, ODE systems are frequently embedded within larger multiphysics simulations [2, 6] and digital twins, and poor calibration at the ODE level propagates upstream, degrading overall fidelity, stability, and decision quality.
@@
A variety of strategies exist to calibrate these ODE models to data. Problem-specific heuristic methods, where techniques such as equation linearization reduce the estimation problem to a linear regression [7, 8, 9], have been demonstrated in several use cases, but suffer from robustness and generalizability issues[10]. Population-based algorithms (like genetic algorithms or swarm optimization) have been successfully used in applications such as battery modeling[11, 12], structural engineering [13], and piezoelectrics [14]. However, population-based methods suffer from poor convergence properties near local minima, and scale poorly with search space dimension. This has been addressed in the past using gradient-based search methods near local optima. In particular, derivatives obtained using automatic differentiation instead of numerical differences have been shown considerable success in solving ODE and PDE (Partial Differential Equation) model identification problems [15, 16, 17], particularly when the models exhibit stiffness characteristics [10, 18]. By backpropagating through the ODE solve—via algorithmic differentiation of the integrator or adjoint sensitivity methods—one obtains exact derivatives of the time-discretized simulation [19] output with respect to parameters, compared to numerical differencing which have inherent discretization errors.
@@
However, practical adoption of efficient model identification algorithms is hindered by substantial software and workflow barriers. In particular, writing fully differentiable simulation code requires nontrivial fluency in state-of-the-art AD frameworks (e.g., PyTorch, JAX, Julia) and with their distinct programming models. For example, JAX’s functional, transformation-driven style (pure functions with explicit jit/vmap/grad semantics) can be unfamiliar and unforgiving for domain scientists. Beyond framework fluency, practitioners must hand-engineer complete optimization pipelines—defining objectives and constraints, organizing multi-experiment datasets, selecting solvers and tolerances, configuring sensitivity modes, and managing compilation and hardware particulars—each step demanding bespoke code and iterative tuning. These burdens reduce accessibility to modern capabilities such as reverse-mode automatic differentiation, just-in-time (jit) compilation, and advanced optimizers. This motivates tooling that abstracts these concerns and automates routine choices while preserving expert control.
@@
Recent advances in artificial intelligence have delivered substantial improvements in user workflow efficiency, and several studies have investigated using LLMs to improve workflows in scientific computing. Kashefi and Mukerji [20] investigated the use of ChatGPT for writing and debugging numerical algorithms for a variety of complex problems. Dong et al. [21] and Pandey et al. [22] investigated using Large Language Models (LLMs) to automate the setup and running of complex fluid dynamics simulations, reducing the need for domain expertise and leading to smoother workflows. Further work has also explored using autonomous AI Agents to automate and accelerate engineering. Elrefaie et al. [23] propose a multi-agent framework for aerodynamic design, streamlining key steps in the automotive design process such as concept sketching, shape retrieval, meshing and simulation. Picard et al. [24] explored utilizing Vision-Language Models (VLMs) in tasks relevant to engineering design such as similarity analysis, CAD generation, and topology optimization. These works demonstrated the potentially large impact AI agents and workflows could have in improving the efficiency of engineers and designers developing new products, similar to the already-realized productivity gains made in fields like software engineering and data analytics [25].
@@
In this work, the use of agentic AI workflows is investigated to streamline parameter identification for complex ODE systems. The proposed framework abstracts away much of the human effort and multidisciplinary expertise traditionally required to formulate and solve these problems while remaining compatible with modern differentiable programming stacks. By delegating routine orchestration to an AI model and exposing only high-level controls to the user, the proposed workflow lowers the barrier to employing just-in-time compilation, automatic differentiation, and advanced optimization methods for ODE calibration. The result is a more accessible, reproducible, and efficient path from problem statement to calibrated model, without sacrificing numerical rigour or expert oversight. The framework is demonstrated on two common types of parameter estimation problems: recovering the parameters from simulated data, and obtaining models that best describe experimentally obtained data.
@@
    Simplified usage interface: A human-readable XML interface and standard Python functions are used to define the problem. These automatically translate into a well-posed constrained optimization in a few steps; capturing objectives, bounds, and constraints—thereby eliminating boilerplate scripting and reducing specification errors.
@@ 2 Proposed Workflow
A three-step, agentic workflow is presented that turns a lightweight, human-readable specification into a compiled, parallel, and differentiable parameter-estimation pipeline. The aim is to minimise manual setup: users describe the problem once, and the agent creates code skeletons, validates consistency, and executes a robust two-stage optimization—producing fitted parameters and solutions requested by the user. The workflow can be visualized in Figure 1, and is briefly described:
@@
    Step I: The user supplies a human-readable XML spec (states, parameters/bounds, dataset). The agent emits a minimal Python skeleton (ODE, loss, quantities of interest) to fill out, so only domain logic must be filled in.
@@
    Step II: After the user completes the functions, the agent cross-checks XML and code, validates shapes/signatures/bounds, separates critical errors from warnings, and auto-remediates issues in short iterations.
@@
    Step III: The agent converts Python callables to JAX code jit-compiles pure functions, sets parallelization, runs a two-stage optimization (PSO for global search, then gradient refinement), and writes fitted parameters and reports to file.
@@ 2.3 Step III: Compilation, Optimization and Post Processing
    Global Search: A globally explorative particle-swarm optimization (referred to as PSO henceforth) to locate promising regions of the parameter space, parallelized across devices (See Appendix A.1).
@@
    Local Refinement: A gradient-based refinement stage, where derivatives are obtained via automatic differentiation through the ODE solve, enabling stable, high-precision updates. This is done similarly to Neural ODEs, and is referred to as the NODE stage henceforth (See Appendix A.2).
@@ 3 Results
3.1 Fitting simulated data
@@ 3.1.2 Van der Pol Oscillator
3.2 Fitting experimental data
This section explores using experimental data to estimate ODE models. This represents a more difficult class of problems, as the data may be sparse or not representative enough, resulting in an ill-posed estimation problem. In addition, experimental data is often noisy, unevenly sampled, and is not available for all integrable variables in the ODE model. Further, the ODE models being estimated are only approximate representations of the physics, making perfect matches to experiment unlikely. The ODE systems are usually inspired by physically complex, already understood equations, and in many cases are difficult to estimate using experimental data. In subsequent subsections, the framework is demonstrated on such examples inspired by real-world, practically relevant studies.
@@ 4 Discussion and Conclusion
This work’s primary contribution is an agentic AI workflow that turns high-level problem intent into an executable, performant parameter-estimation pipeline with minimal manual plumbing. Given human-readable user inputs, the agent automatically creates Python routine skeletons for users to fill, translates user routines into jit-able kernels, checks user inputs for correctness, and orchestrates a global-then-local optimization. The framework is demonstrated on data obtained from simulations of stiff ODE systems, showing that the generated setup is able to recover the parameters that setup the simulation successfully. The framework was also demonstrated on experimental data from several real-world experiments, and the framework returned approximate ODE model systems that are good representations of the experimental data.
@@
This work is intended to be a foundation in intelligent inverse problem solutions, and can be extended in several ways. PSO can be augmented or replaced with Bayesian optimization to cut simulation budgets and improve global search in stiff, weakly identifiable spaces. The work can be extended to include PDE parameter estimation by including new estimation methods like Physics Informed Neural Networks (PINNs) [34] and setting up differentiable finite element or finite volume simulations. An exciting extension would be extension of the agents to do automatic hyper-parameter and ODE solver setting tuning; this includes solver tolerances, time-stepping methods and stepsize controller settings.

### v3

@@ Abstract
Parameter identification for mechanistic Ordinary Differential Equation (ODE) models underpins prediction and control in several applications, yet remains a manual and labor-intensive process: datasets are noisy and partial, models can be stiff and complex, and differentiable implementations demand framework expertise. An agentic AI workflow is presented that converts a lightweight, human-readable specification into a compiled, parallel, and differentiable model calibration pipeline. Users supply an XML description of the problem and fill in a Python code skeleton; the agent automatically validates consistency between problem definition and code, and auto-corrects pathologies in the input deck. It transforms Python callables into pure JAX functions for efficient just-in-time compilation and parallelization. The system then orchestrates a two-stage search comprising global exploration of the parameter space followed by gradient-based refinement. The result is an AD-native, reproducible workflow that lowers the barrier to advanced calibration while preserving expert control. An open-source implementation with a documented API and examples is released, enabling rapid movement from problem statement to interpretable ODE models with minimal effort.
@@ 1 Introduction
Ordinary differential equation (ODE) parameter fitting is a foundational task for turning mechanistic models or ideas into predictive tools across the sciences and engineering. In battery models, ODEs govern degradation kinetics, thermal behavior, and equivalent-circuit dynamics [1, 2]; accurate parameters are essential for state estimation, control, and safety analysis. In combustion modeling, reactions are often represented using Arrhenius ODE systems [3, 4]. In biology, systems of ODEs describe gene regulation, signaling, and epidemiological dynamics [5]. More broadly, ODE systems are frequently embedded within larger multiphysics simulations [2, 6] and digital twins, and poor calibration at the ODE level propagates downstream, degrading overall fidelity, stability, and decision quality.
@@
A variety of strategies exist to calibrate these ODE models to data. Problem-specific heuristic methods, where techniques such as equation linearization reduce the estimation problem to a linear regression [7, 8, 9], have been demonstrated in several use cases, but suffer from robustness and generalizability issues [10]. Population-based algorithms (like genetic algorithms or swarm optimization) have been successfully used in applications such as battery modeling[11, 12], structural engineering [13], and piezoelectrics [14]. However, population-based methods suffer from poor convergence properties near local minima, and scale poorly with search space dimension. This has been addressed in the past using gradient-based search methods near local optima. In particular, derivatives obtained using automatic differentiation instead of numerical differences have shown considerable success in solving ODE and PDE (Partial Differential Equation) model identification problems [15, 16, 17], particularly when models exhibit stiffness characteristics [10, 18]. By backpropagating through the ODE solve via algorithmic differentiation of the integrator or adjoint sensitivity methods, exact derivatives of the time-discretized simulation [19] output are obtained with respect to parameters, compared to numerical differencing which has inherent discretization errors.
@@
However, practical adoption of efficient model identification algorithms is hindered by substantial software and workflow barriers. In particular, writing fully differentiable simulation code requires nontrivial fluency in state-of-the-art AD frameworks (e.g., PyTorch, JAX, Julia) and with their distinct programming models. For example, JAX’s functional, transformation-driven style (pure functions with explicit jit/vmap/grad semantics) can be unfamiliar for domain scientists. Beyond framework fluency, practitioners must hand-engineer complete optimization pipelines—defining objectives and constraints, organizing multi-experiment datasets, selecting solvers and tolerances, configuring sensitivity modes, and managing compilation and hardware particulars—each step demanding bespoke code and iterative tuning. These burdens reduce accessibility to modern capabilities such as reverse-mode automatic differentiation, just-in-time (jit) compilation, and advanced optimisers. This motivates tooling that abstracts these concerns and automates routine choices while preserving expert control.
@@
Recent advances in artificial intelligence have delivered substantial improvements in user workflow efficiency, and several studies have investigated using LLMs to improve workflows in scientific computing. Kashefi and Mukerji [20] investigated the use of ChatGPT for writing and debugging numerical algorithms for a variety of complex problems. Dong et al. [21] and Pandey et al. [22] investigated using Large Language Models (LLMs) to automate the setup and running of complex fluid dynamics simulations, reducing the need for domain expertise and leading to smoother workflows. Further work has also explored using autonomous AI Agents to automate and accelerate engineering. Elrefaie et al. [23] propose a multi-agent framework for aerodynamic design, streamlining key steps in the automotive design process such as concept sketching, shape retrieval, meshing and simulation. Picard et al. [24] explored using Vision-Language Models (VLMs) in tasks relevant to engineering design such as similarity analysis, CAD generation, and topology optimization. These works demonstrated the potentially large impact AI agents and workflows could have in improving the efficiency of engineers and designers developing new products, similar to the already-realized productivity gains made in fields like software engineering and data analytics [25].
@@
In this work, the use of agentic AI workflows is investigated to streamline parameter identification for complex ODE systems. The proposed framework abstracts away much of the human effort and multidisciplinary expertise traditionally required to formulate and solve these problems while remaining compatible with modern differentiable programming stacks. By delegating routine orchestration to an AI model and exposing only high-level controls to the user, the proposed workflow lowers the barrier to employing just-in-time compilation, automatic differentiation, and advanced optimization methods for ODE calibration. The result is a more accessible, reproducible, and efficient path from problem statement to calibrated model, without sacrificing numerical rigor or expert oversight. The framework is demonstrated on two common types of parameter estimation problems: recovering the parameters from simulated data, and obtaining models that best describe experimentally obtained data.
@@
    Simplified usage interface: A human-readable XML interface and standard Python functions are used to define the problem. These are automatically translated into a well-posed constrained optimization in a few steps; capturing objectives, bounds, and constraints—thereby eliminating boilerplate scripting and reducing specification errors.
@@ 2 Proposed Workflow
A three-step, agentic workflow is presented that turns a lightweight, human-readable specification into a compiled, parallel, and differentiable parameter-estimation pipeline. The aim is to minimize manual setup: users describe the problem once, and the agent creates code skeletons, validates consistency, and executes a robust two-stage optimization—producing estimated parameters and solutions requested by the user. The workflow can be visualized in Figure 1, and is briefly described:
@@
    Step I: The user supplies a human-readable XML spec (states, parameters/bounds, dataset). The agent creats a minimal Python skeleton (ODE, loss, quantities of interest) to fill out, so only domain logic must be filled in.
@@
    Step II: After the user completes the functions, the agent cross-checks XML and code, validates shapes/signatures/bounds, separates critical errors from warnings, and auto-remediates issues in iteratively.
@@
    Step III: The agent converts Python callables to JAX code, jit-compiles pure functions, sets parallelization, runs a two-stage optimization (PSO for global search, then gradient refinement), and writes fitted parameters and reports to file.
@@ 2.3 Step III: Compilation, Optimization and Post Processing
    Global Search: A globally explorative particle-swarm optimization (referred to as PSO henceforth) to locate promising regions of the parameter space, parallelized across devices (See Appendix A.1 for details).
@@
    Local Refinement: A gradient-based refinement stage, where derivatives are obtained via automatic differentiation through the ODE solve, enabling stable, high-precision updates. This is done similarly to Neural ODEs, and is referred to as the NODE stage henceforth (See Appendix A.2 for details).
@@ 3 Results
3.1 Estimating parameters from simulated data
@@ 3.1.2 Van der Pol Oscillator
3.2 Estimating parameters from experimental data
This section explores using experimental data to estimate ODE models. This represents a more difficult class of problems, as the data may be sparse or not representative to accurately capture all phenomena of interest, resulting in an ill-posed estimation problem. In addition, experimental data is often noisy, unevenly sampled, and is not available for all integrable variables in the ODE model. Further, the ODE models being estimated are only approximate representations of the physics, making perfect matches to experiment unlikely. The ODE systems are usually inspired by physically complex, already understood equations, and in many cases are difficult to estimate using experimental data. In subsequent subsections, the framework is demonstrated on such examples inspired by real-world, practically relevant studies.
@@ 4 Discussion and Conclusion
This work’s primary contribution is an agentic AI workflow that turns high-level problem intent into an executable, performant parameter-estimation pipeline with minimal user effort. Given human-readable user inputs, the agent automatically creates Python routine skeletons for users to fill, translates user routines into jit-able kernels, checks user inputs for correctness, and orchestrates a global-then-local optimization. The framework is demonstrated on data obtained from simulations of stiff ODE systems, showing that the generated setup is able to recover the parameters that setup the simulation successfully. The framework was also demonstrated on experimental data from several real-world experiments, and the framework returned approximate ODE model systems that are good representations of the experimental data.
@@
This work is intended to be a foundation in intelligent inverse problem solutions, and can be extended in several ways. PSO can be augmented or replaced with Bayesian optimization to cut simulation budgets and improve global search in stiff, weakly identifiable spaces. The work can be extended to include PDE parameter estimation by including new estimation methods like Physics Informed Neural Networks (PINNs) [34] and setting up differentiable finite element or finite volume simulations using frameworks such as PhiFlow [35]. Further extension of the agent can be done to automate hyper-parameter tuning and ODE solver settings; this includes solver tolerances, time-stepping methods and stepsize controller settings.